OpenStudy (jazzyfa30):

Find the value of the discriminant and describe the number and type of roots. 3x2 - 8x + 2 = 0

OpenStudy (jazzyfa30):

@tkhunny

OpenStudy (tkhunny):

\(b^{2} - 4ac\)? This should ring bells. Have you used the quadratic formula?

OpenStudy (jazzyfa30):

no this is new to me

OpenStudy (tkhunny):

You've seriously never used the quadratic formula? Just checking.

OpenStudy (jazzyfa30):

no never

OpenStudy (tkhunny):

(sigh) Why are you given problem that you have no tools to solve? Have you ever solved a quadratic equation by "Completing the Square"?

OpenStudy (jazzyfa30):

no

OpenStudy (tkhunny):

Gaaa!! Who writes these course materials?! Okay, a GENERAL quadratic in STARNDARD form is this \(ax^{2} + bx + c = 0\), where a, b, and c are Real Numbers. Do you believe?

OpenStudy (jazzyfa30):

yes

OpenStudy (tkhunny):

One day, you should learn the process called "Completing the Square". I guess we'll have to wait for another lesson on another day for that. Suffice it to say that if you use this technique to SOLVE this quadratic equation, you will get: \(x = \dfrac{-b \pm \sqrt{b^{2} - 4ac}}{2a}\) Since we can't help you prove that, as you haven't the tools, you'll jus have to believe me. a, b, and c in the formula are the same as we just saw in the Standard Form. Following?

OpenStudy (jazzyfa30):

yes

OpenStudy (tkhunny):

This will sound like a silly question. What value for 'b' will make this formula fail?

OpenStudy (jazzyfa30):

idk

OpenStudy (tkhunny):

Actually, that's an excellent answer. There is no such value. 'b' is just 'b'. It tells us nothing by itself. Okay, what value for 'a' can make this formula fail? Think carefully! "idk" is not acceptable on this one.

OpenStudy (jazzyfa30):

nothing

OpenStudy (tkhunny):

?? If by "nothing" you mean a = 0 is bad, then you are exactly correct. You can't have zero (0) in the denominator, can you?

OpenStudy (jazzyfa30):

no you cant have 0 in a denominator

OpenStudy (tkhunny):

Great, so a better not be zero. A review of the standard form ax^2 + bx + c = 0 suggests for a = 0, it isn't even a quadratic equation! So, a better not be zero - ever! One more silly question. What sorts of values can you put in a square root? \(\sqrt{4}\) -- Does that work? \(\sqrt{0}\) -- Does that work? \(\sqrt{-3}\) -- Does that work?

OpenStudy (jazzyfa30):

4 yes but the rest is no cus nothing multiplies by itself to get that particular number such as 0 and -3

OpenStudy (tkhunny):

Careful. 0*0 = 0. That one works. What we are saying, for now, is that a square root can handle anything that is NOT NEGATIVE. Agreed?

OpenStudy (jazzyfa30):

yes

OpenStudy (tkhunny):

Okay, we are almost there. Tired yet? Way back at the beginning, I wrote the quadratic formula. We talked about the 'b' out in front. We talked about the 'a' in the denominator. The only part we have not talked about is the square root. \(\sqrt{b^{2} - 4ac}\) Since we just decided that square roots can't handle anything negative, if we get \(b^{2} - 4ac < 0\) we are just totally out of luck if we want to find a Real Number. Does this make sense?

OpenStudy (jazzyfa30):

yes kinda

OpenStudy (tkhunny):

Suppose we have \(2x^{2} + 3x + 4 = 0\) a = 2 -- We are very happy this is not zero. b = 3 c = 4 \(b^{2} - 4ac = 3^{2} - 4(2)(4) = 9 - 32 = -23\) If we are to find a solution, we will need to evaluate \(\sqrt{-23}\). Since we don't like \(\sqrt{-23}\), we throw our hands in the air and exclaim, "There is no solution in the Real Numbers!" How are we doing?

OpenStudy (jazzyfa30):

great im following

OpenStudy (tkhunny):

Well, that's what happens when \(b^{2} - 4ac < 0\) What do you suppose we get to think if \(b^{2} - 4ac > 0\)?

OpenStudy (jazzyfa30):

9-32>0 +32 +32 9>32 -9 \[\sqrt{23}\] yes i think

OpenStudy (tkhunny):

?? I don't know what you did? 9 - 32 will always be -23. There is nothing you can do to change that. Suppose we have \(2x^{2} +3x - 4 = 0\) a = 2 -- We are very happy this is not zero. b = 3 c = -4 \(b^{2} −4ac=3^{2} −4(2)(-4)=9+32= 41 \) If we are to find a solution, we will need to evaluate \(\sqrt{41}\). Can we do that?

OpenStudy (jazzyfa30):

ohhh i see what i did wrong

OpenStudy (tkhunny):

Good. How about \(\sqrt{41}\)? Is that something that exists? Are we happy with that result?

OpenStudy (jazzyfa30):

yes

OpenStudy (tkhunny):

Perfect. Now, this one has a special case. 41 is pretty ugly sitting under a square root. What do you suppose we can say if \(b^{2} - 4ac\) is a Perfect Square? Side Question: Have you solved a quadratic equation by FACTORING?

OpenStudy (jazzyfa30):

yes

OpenStudy (jazzyfa30):

idk

OpenStudy (tkhunny):

That's good. We'll need that. Take a look at that Quadratic Formula. If \(b^{2} - 4ac\) is a perfect square, we can just plain not worry about the square root. Suppose we have \(x^{2} + 3x − 4 = 0\) a = 1 -- We are very happy this is not zero. b = 3 c = -4 \(b^{2} − 4ac =3^{2} − 4(1)(−4) = 9 + 16 = 25\) The quadratic formula loses quite a bit of its glory. \(x = \dfrac{-3 \pm \sqrt{25}}{2(1)} = \dfrac{-3 \pm 5}{2}\) It's a WHOLE LOT EASIER all of a sudden, isn't it?!

OpenStudy (jazzyfa30):

yes

OpenStudy (tkhunny):

This one has an interesting implication. Without the complication of a square root in the final answer, you probably should feel a little bad about not factoring it. Given a little more effort, factoring can be found. Okay, we're down to the last conversation. What if \(b^{2} - 4ac = 0\)?? What can we tell from that?

OpenStudy (jazzyfa30):

idk

OpenStudy (tkhunny):

Come on. I'm going to make you tell me. Look at the Quadratic formula and imaging what happens if you just throw out the square root part. \(\sqrt{0} = 0\), right?

OpenStudy (jazzyfa30):

ummmm......................................

OpenStudy (tkhunny):

Think about the numerator -b + 0 -b - 0 What is so special about that?

OpenStudy (jazzyfa30):

b is negative

OpenStudy (tkhunny):

?? That makes no sense. Look at the three examples I provided above. b = 3. What is so special about the relationship between -3 + 0 and -3 - 0 ??

OpenStudy (jazzyfa30):

it still gonna be a ngative with both equations

OpenStudy (tkhunny):

Okay, if b = -5 What is special about the relationship between these two expressions? -(-5) + 0 = 5+0 and -(-5) - 0 = 5-0 Look at what changes in the two expressions. What changes and what does it do? Search your mind, Drag some sense out of the relationship!!

OpenStudy (jazzyfa30):

the addition and subtraction symbol

OpenStudy (tkhunny):

Okay, that's what changes. Does it make any difference if we are adding or subtracting zero (0)?

OpenStudy (jazzyfa30):

yes it changes it from being ngative/positive

OpenStudy (tkhunny):

It changes WHAT from being positive or negative? -5 + 0 = ?? -5 - 0 = ??

OpenStudy (jazzyfa30):

the answer

OpenStudy (tkhunny):

Answer those last two questions. Fill in the two "??".

OpenStudy (jazzyfa30):

-5+0=-5 -5-0=5

OpenStudy (tkhunny):

Rethink!!! -5-0 = -5 How did you get +5? If you can't get the arithmetic, you can't do the algebra. Please be more careful.

OpenStudy (tkhunny):

Note: You have been working very hard. You have been very patient. Hang in there just a little longer. You'll get it!

OpenStudy (jazzyfa30):

ok

OpenStudy (tkhunny):

Do you see that, now? -5 + 0 = -5 -5 - 0 = -5 The are EXACTLY THE SAME!

OpenStudy (tkhunny):

Zero doesn't do anything in addition or subtraction.

OpenStudy (jazzyfa30):

yes i see and ok

OpenStudy (jazzyfa30):

im sorry i gtg will u b on 2morrow

OpenStudy (tkhunny):

Okay, this is a problem about the "discriminant". The "discriminant" is that thing that gives us information without having to solve the whole problem. In this case, the "discriminant" is just that \(b^{2} - 4ac\) that we have been beating to death for the last couple hours. Recap of the discriminant Negative? No Real Number Solution Positive? Real Number Solutions Perfect Square? Might have been able to factor it if we had tried harder. Zero? Think to yourself, "I am a pansy, I failed to recognize a perfect square trinomial" Whew!! You probably know more about the discriminant and the quadratic formula than anyone else in your class or study cohort. Are we ready to solve the original problem? Do we remember what it was? Find the value of the discriminant and describe the number and type of roots. 3x2 - 8x + 2 = 0 Go! Let's see what you get.

OpenStudy (tkhunny):

I'll be waiting for your answer.

OpenStudy (jazzyfa30):

@nincompoop

OpenStudy (jazzyfa30):

@mathslover

mathslover (mathslover):

Jazzy, @tkhunny explained well . Where are you having problem?

OpenStudy (jazzyfa30):

i dont understand how to do it

OpenStudy (jazzyfa30):

@tkhunny idk i cant remember